OpenStudy (abbycross167):

Can someone please check my work on this math question? please? I'll fan and give a medal!!

OpenStudy (abbycross167):

Please help me solve this equation – -4.9t2 + 24.5t + 117.6 = 0

OpenStudy (abbycross167):

@mjdennis can you please try to help me?

geerky42 (geerky42):

Have you tried using quadratic formula?

OpenStudy (abbycross167):

no sir... i cant figure out how to use it

geerky42 (geerky42):

Quadratic Formula: \(ax^2+bx+c=0\Longrightarrow x=\dfrac{\text-b\pm\sqrt{b^2-4ac}}{2a}\) Here, you have \(-4.9t^2 + 24.5t + 117.6=0\) So that would make \(a=-4.9,~b=24.5,~c=117.6\) Now plug in values and use calculator to get the answer. Is that clear?

OpenStudy (abbycross167):

ok so -24.5+/-\/([24.5^2]-4*-4.9*117.6 x=------------------------------- 2(-4.9)

OpenStudy (abbycross167):

@ganeshie8 can you try to help me please?

OpenStudy (jianenriquez):

have you tried solving what you just posted above?

OpenStudy (abbycross167):

so so... I can't figure it out

OpenStudy (abbycross167):

hold on, I think i got it

jimthompson5910 (jim_thompson5910):

Let's focus on the part in red first. \[\Large x=\frac{-b \pm \sqrt{{\color{red}{b^2-4ac}}}}{2a}\] This is the discriminant So we have \(\Large b^2-4ac\). Let's plug in the values for a,b,c. a = -4.9 b = 24.5 c = 117.6 \[\Large b^2-4ac = (24.5)^2 - 4*(-4.9)*(117.6) = \underline{ \ \ \ \ \ \ \ \ \ \ \ \ }\] Fill in the blank. Tell me what you get.

OpenStudy (jianenriquez):

Heres the definition of the quadratic formula :) i think you just got confused with the +- Btw, a little piece of advice, never forget b^2 - 4ac

OpenStudy (abbycross167):

so would t=−2.9999999999999996,8 ?

jimthompson5910 (jim_thompson5910):

you should get t = -3 or t = 8

jimthompson5910 (jim_thompson5910):

so it looks like you pretty much got it. You just need to round -2.9999999999999996 to -3

OpenStudy (jianenriquez):

yep, t= -3 and t=8

OpenStudy (abbycross167):

oh right! I always forget to round! thank you!!

jimthompson5910 (jim_thompson5910):

If t is a time value, then only t = 8 makes sense. A negative time value makes no sense.

OpenStudy (abbycross167):

okay thank you

OpenStudy (mjdennis):

Good work. Thanks for asking for me. I forgot to close my OpenStudy window last night, so didn't know to come and help!